Riverbed Corporation has the following four items in its ending inventory:

Item Cost Estimated
Selling Price
Estimated
Disposal Costs
Neutrinos $ 1,950 $2,230 $99
Ocillinos 5,060 4,960 112
Electrons 4,280 4,615 219
Protons 3,420 4,430 91

(a)

Assume that Riverbed is a public company using IFRS. Determine the total value of ending inventory, using the lower of cost and net realizable value model applied on an individual item basis.
Answer #1
Item Cost Net realizable value [Estimated
selling price - Estimated
disposal costs]
Lower of cost and net realizable value
Neutrinos 1,950 (2,230 - 99) = 2,131 1,950
Ocillinos 5,060 (4,960 - 112) = 4,848 4,848
Electrons 4,280 (4,615 - 219) = 4,396 4,280
Protons 3,420 (4,430 - 91) = 4,331 3,420
Total value of ending inventory $14,498